OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write the solution as an ordered pair.

OpenStudy (anonymous):

SAIFOOOOOOOOO!!!!!!!!!!!!!

OpenStudy (saifoo.khan):

VEROOOOOOOOOOO!

OpenStudy (anonymous):

i got it (the answer)! and i hate being called that :/

OpenStudy (saifoo.khan):

In the first scenario, m+5f=25 And for the 2nd, 10f=40 f=4 Substitute f=4 into the 1st scenario m+5*4=25 m+20=25 m=5 therefore, the membership fee is $5 and the session fee is $4

hero (hero):

I thought she wanted to be called Vera

OpenStudy (anonymous):

no. T-O-R-Y!!!!!!!!!!!

hero (hero):

Wait....I'm confused....Isn't your name Veronica?

OpenStudy (anonymous):

ya

hero (hero):

Okay.....

OpenStudy (anonymous):

I'm called Tory for 3 reasons: 1. My last name is Torrente 2. I look like a Victoria 3. Veronica means victory bringer :)

OpenStudy (saifoo.khan):

LOL.

hero (hero):

Okay, but Vera is better.....because it's short for Aloe Vera.....something soothing, soft and smooth....

OpenStudy (anonymous):

Saifoo i guess you have a typing mistake its not 4 its 3